# ReNF: Rethinking the Design Space of Neural Long-Term Time Series Forecasters

### v4

@@ Abstract
Neural Forecasters (NFs) are a cornerstone of Long-term Time Series Forecasting (LTSF). However, progress has been hampered by an overemphasis on architectural complexity at the expense of fundamental forecasting principles. In this work, we return to first principles to redesign the LTSF paradigm. We begin by introducing a Multiple Neural Forecasting Proposition that provides a theoretical motivation for our approach. We propose Boosted Direct Output (BDO), a novel forecasting paradigm that synergistically hybridizes the causal nature of Auto-Regressive (AR) models with the stability of Direct Output (DO). In addition, we stabilize the learning process by smoothly tracking the model’s parameters. Extensive experiments show that these principled improvements enable a simple MLP to achieve state-of-the-art performance, outperforming recent, complex models in nearly all cases, without any specific considerations in the area. Finally, we empirically verify our proposition, establishing a dynamic performance bound and identifying promising directions for future research. The code for review is available at: .
@@ 1 Introduction
The progression of any real-world event is a unique, non-repeatable process, often governed by chaotic dynamics we cannot perfectly measure (robertson1929uncertainty) or describe. This implies that any observed time series is just one stochastic sample from a complex underlying system. Consequently, a fundamental open question persists: how can we best estimate the long-term future states depending only on a single, observed historical sequence?
@@
Deep Neural Networks (DNNs) have recently received increasing attention in Long-Term Time Series Forecasting (LTSF) (survey2025). Their ability to model high-dimensional, non-linear dependencies makes the Neural Forecaster (NF) a promising tool for capturing complex temporal dynamics and implicit dependencies (selfcontrastivenonlinear). However, the literature reveals several critical problems that should be addressed to unlock the full potential of DNNs in this domain.
@@
One confusing puzzle is that both advanced architectures, like transformer-based (patchtst) and simple linear-based models (dlinear), are reported to achieve state-of-the-art performance interchangeably while their complexity varies significantly. This is partly because some datasets prefer parsimonious (deng2024parsimony) while the others have sufficient volume for complex models to fit. But perhaps a deeper issue is the insufficient exploration of the networks’ intrinsic capabilities. As indicated by (lu2025timecapsule), many NFs contain redundant components and are used selectively depending on the property of the data, resulting in their full potential remaining untapped without extensive tuning or regularizations.
@@
Furthermore, the field’s focus has shifted towards designing specialized modules for properties that are believed to be beneficial for LTSF, such as multi-scale (wang2024timemixer) and non-stationary (nonstarionarytransformer). However, the progress has become erratic because these architectural additions often yield subtle gains while overlooking fundamental principles. We re-emphasize that progress in this field becomes circuitous when we introduce advanced models merely according to the empirical conclusions from other realms, but overlook the detailed instructions for their full utilization. A primary and more direct path to advancement may lie in fundamentally improving the training stability and generalization capabilities of NFs themselves.
@@
Beyond model architecture, we find that the dominant Direct Output (DO) framework does not make full use of the available supervision. In a standard DO setup, an NF is trained to predict the entire future horizon in a single forward pass, meaning the label information is leveraged only once per optimization step (see Sec.2.3 for the detailed explanation). This encourages the model to finally learn a monolithic representation that maps the entire history to the entire future, without explicitly modeling the sequential dependencies within the forecast itself. We contend that this approach hinders the model from developing a more granular, causal understanding of the future, thereby limiting its full potential.
@@
To address these fundamental shortcomings, this paper proposes a new LTSF framework, derived from first principles, that establishes a reliable and high-performing benchmark model. Our contributions are summarized as follows:
  • We propose the Multiple Neural Forecasting Proposition (MNFP) with empirical evaluations, which provides a formal motivation for employing multiple Neural Forecasters (NFs) in Long-Term Time Series Forecasting (LTSF).
@@
  • We redesign the forecasting architecture by introducing a novel, streamlined forecasting paradigm that synergistically combines the strengths of Direct Output (DO) and Auto Regressive (AR) methods. Furthermore, we apply the Exponential Moving Average (EMA) technique to effectively stabilize the convergence of Neural Forecasters.
@@
  • We conduct extensive experiments to demonstrate that a pure MLP-based forecaster, when trained with our paradigm, can significantly outperform recent state-of-the-art models on nearly all standard LTSF benchmarks. These results establish a new and efficient performance baseline for the field.
@@ 2 Method
In this section, we lay out the theoretical and methodological foundations of our work. We begin by introducing a core proposition that provides the theoretical underpinnings for our approach. To build intuition, we then illustrate our designs using a simple MLP architecture on carefully selected datasets that highlight the key effects of our methods. The comprehensive experimental setup and full results on benchmark datasets are detailed in Sec.3 and the Appendix.
2.1 Preliminaries
Definition 1 (Neural Forecasting Machine).
A Forecasting Machine (FM) is an end-to-end model described by the four-tuple $(t_{x},t_{y},\theta,\gamma)$, where $t_{x}$ and $t_{y}$ denote the input and output lengths. The symbol $\theta$ denotes the model’s parameters, and $\gamma$ denotes the random state encompassing all random factors such as the computational environment and exogenous variables.
We define the NFM as an FM modeled by neural nets and satisfying that, given an input time series $X\in\mathbb{R}^{t_{x}}$, an NFM with a determined property tuple $(t_{x},t_{y},\theta,\gamma)$ yields one and only one series $Y\in\mathbb{R}^{t_{y}}$ through the operation $Y=\text{NFM}[t_{x},t_{y},\theta,\gamma](X)$.
Definition 2 (Forecasting Task).
We represent a time series forecasting task with the tuple ($X_{h},X_{f},\hat{Y}_{f},Y_{f}$), where $X_{h},X_{f}$ denote the history and future portions of an observed time series. $Y_{f}$ denotes the expected forecast of $X_{f}$, while $\hat{Y}_{f}$ is an empirical forecast generated by a model such as an NFM. The goal of a forecasting task is to make $\hat{Y}_{f}$ as close as possible to $X_{f}$ with the input $X_{h}$.
@@ 2.2 Theoretical Motivation for Multiple Forecasts
The following proposition provides the theoretical motivation for our work. For simplicity, we just present the proof for the univariate case; however, the proposition can be readily extended to the multivariate setting.
Figure 1: Illustration of the MNFP.
Proposition 1 (Multiple Neural Forecasting Proposition (MNFP)).
@@
Given a NFM $\Phi(t_{x},t_{y},\theta,\gamma)$ and an observed time series $X_{h}=(x_{1},x_{2},\cdots,x_{n})$ where each element $x_{t}$ is drawn from a true distribution $p_{t}(\mu_{t},\sigma_{t}^{2})$ with mean $\mu$ and standard deviation $\sigma_{t}$. One can generate a series by $\Phi$: $\hat{Y}_{f}=(y_{1},y_{2},\cdots,y_{T})$ with $y_{t}$ drawn from the expected forecast distribution $\hat{p}(\hat{\mu_{t}},\hat{\sigma}^{2}_{t})$ such that $\hat{Y}_{f}$ approaches $Y_{f}$ almost surely, and the error between $\hat{Y}_{f}$ and $X_{f}$ is bounded.
Specifically, a trivial but informative upper bound on the $l_{2}$ error is given by $T(\lambda+\sqrt{c}(b+\sigma_{t}))/\sqrt{c}$ under the following mild conditions:
  • (A1) The elements in the time series are bounded by a positive real constant $\lambda$.
  • (A2) The NFM generates candidate series $\{\hat{Y}^{(i)}_{f}\}_{i=1}^{c>1}$ under various $\theta^{i}$ or random states $\gamma^{i}$.
  • (A3) The predictive bias of the NFM is bounded: $\sup_{t}{|{\hat{\mu}}_{t}-\mu_{t}|}=b$
The proof and further analysis can be found in Appendix C. To interpret, the shown bound value provides us with some intuitions: for instance, the forecasting error of the NFM naturally grows with the forecast horizon $T$ and the data range $\lambda$, because of the expanded solution space. We emphasize that condition A3 distinguishes between the observed real series and the expected true series. This frames the observed data not as the absolute ground truth, but as a single, near-optimal sample from the underlying stochastic process. While it is feasible to theoretically approach the expected forecast, the empirical gap between any observation $x_{t}$ and the true expectation $\mu_{t}$ is practically irreducible.
The most critical insight of the proposition is that an accurate forecast is theoretically achievable even with a rather weak generator, provided a sufficiently large number of candidate outputs is produced. This proposition is rooted in the well-established theory of ensemble methods (bates1969combination), which demonstrates that combining multiple diverse estimators can lead to a more accurate and robust final estimate. Furthermore, higher-quality historical data and a more powerful NFM (which reduces the estimation error $b$) both serve to constrain the solution space, leading to better predictions. Inspired by this, we can propose the following conception,
@@
Post combination of multiple forecasts. Given a set of candidate outputs from an NFM, there exist combinatorial methods to synthesize a single, integrated forecast that would be more accurate than any individual candidate. We can represent such a method as an unknown functional $g_{c}$ that produces the post-combined forecast as:
@@
In our experiments, we will empirically show the potential of this combination on real-world datasets. In practice, however, directly creating an optimal combination function without access to future information is a significant challenge (combiningreview). Therefore, designing a framework that can implicitly and effectively leverage these multiple forecasts is one of the key principles guiding our subsequent work.
@@ 2.3 Novel Forecasting Paradigm
The current frameworks in LTSF involve two main strategies: 1-step Auto Regressive (AR(1)) and Direct Output (DO). While AR models perform consistently with recurrent state space models like RNNs (lstm4ltsf), they are known to suffer from significant error accumulation in long-horizon forecasting and are empirically evidenced to largely fall behind DO in LTSF (zhou2021informer), making a strong reason for the dominant position of DO in LTSF. However, the shortcomings of the DO approach itself have been largely overlooked. We argue that despite its simplicity, the most significant weakness of DO, particularly when compared to AR, is its lack of inherent causality.
To illustrate this, we conduct a simple experiment in Fig.3: we consecutively split a forecast of length $L$ into $m$ non-overlapped sub-forecasts with length $L/m$, and each part is predicted by an independent linear head from a shared representation. The final forecast is a concatenation of these segments. Empirically, it turns out that the performance of this multi-headed model is nearly identical to a standard DO model with a single head predicting the full horizon. This result suggests that the NF is unaware of the temporal relationships within the future sequence and makes the prediction without considering the distance/interval between the input history and output future. As a result, it learns to forecast each segment independently, which is fundamentally anti-intuitive.
Figure 3: We make the forecast by applying independent heads on several non-overlapped chunks. The right figure shows the learning process in different settings.
According to this observation, we propose a new forecasting paradigm as follows,
Definition 3 (Boost Direct Output (BDO)).
@@
Intuitively, BDO recursively generates forecasts for progressively longer horizons, reusing predictions from previous stages. This incorporates an AR-like causal structure into the forecasting process, while retaining the patch-wise output characteristic of DO that mitigates significant error accumulation. A balance between these two properties can be achieved by properly setting the number of recursive stages, $N$. Furthermore, since short-term forecasting is generally an easier task than long-term forecasting, BDO effectively creates a learning curriculum. We can reasonably expect this paradigm to outperform both pure AR and pure DO strategies, especially in the challenging LTSF setting.
@@ 2.4 Model Architecture
Figure 4: Model Structure of ReNF.
We construct our NFs using only MLP and linear layers for three primary reasons: 1) As foundational deep learning modules, improvements demonstrated on them are broadly applicable and convincing. 2) Their computational efficiency (low FLOPs) facilitates rapid experimentation and verification. 3) Recent work has shown that simple MLP-based architectures are often sufficient for a wide range of forecasting tasks (ekambaram2023tsmixer).
In principle, we build two NF variants, ReNF-$\alpha$ and ReNF-$\beta$, which differ in their degree of non-linearity to better handle datasets of varying complexity. The overall architecture is shown in Fig.4. We stack multiple blocks to adopt the BDO strategy. Each consists of a linear layer for representation projection, followed by an MLP for nonlinear transformation. Each block is also equipped with a dedicated linear head that maps the internal representation back to the data space, steering it to function as a sub-forecaster for horizon $h_{j},j=0,1,...,k$. In the recursive BDO process, the output of each sub-forecaster is concatenated with the original input data to form the new input for the next one. To allow for deeper representation flow, the ReNF-$\beta$ variant also incorporates skip-connections between the representation spaces of consecutive blocks.
We employ RevIN (kim2021reversible) as the pre-normalization for the initial input data to reduce the distribution discrepancy between the training and evaluation phases. For consistency, we also apply pre-normalizations to the input of each sub-forecaster. Additionally, we apply dropout before the initial linear projection to prevent the model from being overly dependent on the history observation and the concatenated information in the data space.

### v5

# ReNF: Rethinking the Principles of Neural Long-Term Time Series Forecasters

@@ Abstract
Neural Forecasters (NFs) have become a cornerstone of Long-term Time Series Forecasting (LTSF). However, recent progress has been hampered by an overemphasis on architectural complexity at the expense of fundamental forecasting principles. In this work, we revisit the principles of LTSF. We begin by formulating a Variance Reduction Hypothesis (VRH), positing that generating and combining multiple forecasts is essential to reducing the inherent uncertainty of NFs. Guided by this, we propose Boosted Direct Output (BDO), a streamlined paradigm that synergistically hybridizes the causal structure of Auto-Regressive (AR) with the stability of Direct Output (DO), while implicitly realizing the principle of forecast combination within a single network. Furthermore, we address the critical validation-test generalization gap by employing parameter smoothing to stabilize optimization. Extensive experiments demonstrate that these trivial yet principled improvements enable a direct temporal MLP to outperform recent, complex state-of-the-art models in nearly all benchmarks, without relying on intricate inductive biases. Finally, we empirically verify our hypothesis, establishing a dynamic performance bound that highlights promising directions for future research. The code for review is available at: https://anonymous.4open.science/r/ReNF-A151.
@@ 1 Introduction
The progression of any real-world event is a unique, non-repeatable process, often governed by chaotic dynamics that we cannot perfectly measure (robertson1929uncertainty) or describe. This implies that any observed time series is a single stochastic realization of a complex underlying system, inevitably corrupted by measurement errors. Consequently, a fundamental open question persists: how can we best estimate long-term future states using a data-driven model that relies solely on a single observed historical sequence?
@@
Deep Neural Networks (DNNs) have recently received considerable critical attention in Long-Term Time Series Forecasting (LTSF) (survey2025). Their capacity to model high-dimensional, non-linear dependencies makes the Neural Forecaster (NF) a promising tool for capturing complex temporal dynamics (selfcontrastivenonlinear). However, the literature reveals several critical bottlenecks that have to be addressed to unlock the full potential of DNNs in this domain.
@@
A confounding paradox in recent research is that advanced architectures, such as Transformer-based models (patchtst), and simple linear models (dlinear) are reported to achieve state-of-the-art performance interchangeably, despite vast differences in complexity. This is partly because some datasets favor parsimonious models (deng2024parsimony), while others possess sufficient volume to fit complex architectures. Yet, a deeper issue may be the insufficient exploration of the networks’ intrinsic capabilities. As indicated by (lu2025timecapsule), many complex NFs contain redundant components that are activated selectively based on data properties, leaving their full potential untapped without extensive tuning.
@@
Furthermore, the field has shifted towards designing specialized modules for specific properties, such as multi-scale modeling (wang2024timemixer) and non-stationarity adaptation (nonstarionarytransformer). However, progress has become erratic, as these architectural additions often yield subtle gains while overlooking fundamental principles. A primary and more general path to advancement may lie in fundamentally improving the training stability and generalization capabilities of NFs themselves.
Notably, training on time series often suffers from significant instability. The non-stationary nature of real-world time series, combined with chronological data splitting, leads to both internal (batch-to-batch) and external (train/validation/test) distribution mismatches. This causes the optimization path to be heterogeneous, rendering the learning process for NFs ineffective. Specifically, internal data redundancy can cause the model to overfit spurious patterns (liu2024timebridge), while external distribution shift creates inconsistencies across training phases. Although normalization techniques (kim2021reversible) address train-test shifts, a critical discrepancy remains between validation and testing performance, which insidiously prevents NFs from demonstrating their true capabilities.
@@
Beyond model architecture, we identify a structural limitation in the dominant Direct Output (DO) framework: it fails to fully utilize available supervision. In a standard DO setup, an NF is trained as a single-input single-output system to predict the entire future horizon in a single forward pass, meaning the whole label information is leveraged only once per optimization step. This may tend to encourage the model to learn a monolithic structural representation that maps history directly to the future, without explicitly modeling the sequential dependencies within the forecast itself. We contend that this approach hinders the model from developing a more granular, causal understanding of the future, thereby limiting its full potential.
@@
To address these fundamental issues, this paper proposes some principled improvements that help establish a reliable and high-performing NF. Our contributions are summarized as follows:
@@
  • We redesign the forecasting architecture with a novel, streamlined stacking paradigm that synergistically combines the causal structure of Auto-Regressive (AR) methods with the stability of Direct Output (DO).
  • We identify the validation-test generalization gap as a critical bottleneck in evaluating Neural Forecasters. To bridge this gap, we adopt Exponential Moving Average (EMA), demonstrating that this standard technique is particularly effective in the LTSF domain for stabilizing optimization and preventing overfitting.
@@
  • We conduct extensive experiments to evaluate the effectiveness of our proposals and demonstrate that a pure MLP-based forecaster, when trained with our paradigm, can still outperform recent state-of-the-art models on nearly all standard LTSF benchmarks, establishing a new and efficient performance baseline for the field.
@@ 2 Method
In this section, we lay out the methodological foundations of our work. We begin by introducing a core hypothesis that provides the primary motivation for our multi-forecasts approach. To build intuition, we highlight the key effects of each consideration with empirical demonstrations using a toy MLP on selected datasets.
Problem Formulation. We define a Neural Forecasting Machine (NFM) as a mapping $\Phi_{\theta}(X,\gamma)\to Y$, where $\theta$ denotes learnable parameters and $\gamma$ represents the model’s stochastic state (e.g., random seeds). Given a dataset of finite realizations drawn from an underlying process, partitioned into history $X_{h}\in\mathbb{R}^{T_{x}}$ and future $X_{f}\in\mathbb{R}^{T_{y}}$, the forecasting task aims to learn a NFM such that the generated prediction $\hat{Y}_{f}=\Phi_{\theta}(X_{h},\gamma)$ minimizes the error metric between $\hat{Y}_{f}$ and $X_{f}$.
@@ 2.1 Theoretical Motivation for Multiple Forecasts
We allow the trained NFM as a generator capable of producing multiple outputs for a fixed input, conditioned on varying internal states or configurations. To theoretically assess the capability of such a forecaster, we estimate the point-wise error between the generated predictions $\hat{Y}_{f}$ and the expected future $X_{f}$ as follows,
Figure 1: A single trained NFM $\Phi_{\theta}$ can generate multiple forecasts from a fixed input $X_{h}$ under various states $\gamma$ and $\theta$. These forecasts are expected to follow the empirical process $(\hat{p}_{t},~t=1,2,\cdots)$ approximated by the NFM with the observed data. The model-estimated process is distinguished from the real data distribution $(p_{t},~t=1,2,\cdots)$ by a bias.
Hypothesis 1 (Variance Reduction Hypothesis (VRH)).
@@
Given a NFM $\Phi_{\theta}$ and an observed time series $X_{h}=(x_{1},x_{2},\cdots,x_{h})$ where each element $x_{t}$ is drawn from the true distribution $p_{t}(\mu_{t},\sigma_{t}^{2})$ with mean $\mu$ and standard deviation $\sigma_{t}$. We hypothesis that one can generate a series from the forecasts of $\Phi$: $\hat{Y}_{f}=(y_{1},y_{2},\cdots,y_{N})$ with each element $y_{t}$ drawn from the estimated future distribution $\hat{p}(\hat{\mu_{t}},\hat{\sigma}^{2}_{t})$ such that the absolute error between $\hat{Y}_{f}$ and $X_{f}$ relates to a trivial upper bound,
$(\underbrace{\sigma_{t}}_{\textbf{data~variance}}+\underbrace{b}_{\textbf{predictive~bias}}+\underbrace{\lambda/\sqrt{c}}_{\textbf{predictive~variance}})*N$ (1)
where $b:=\sup_{t}{|{\hat{\mu}}_{t}-\mu_{t}|}$, $\lambda$ can be the upper bound of the second moment of any single forecast, and $c$ denotes the number of series.
Remark. This bound relies on independence assumptions, which simplifies the dependencies inherent in our empirical design. Nonetheless, the analysis provides a qualitative guide, exposing the problem of predictive variance and motivating the shift from single-shot to multi-output forecasters.
The detailed proof is provided in Appendix B.1. To interpret this bound value, we decompose the error into three distinct components. First, $\sigma_{t}$ represents the intrinsic system uncertainty, which is irreducible by the model. Second, the predictive bias reflects the approximation gap between the learned and true distributions; while this can be reduced by enhancing model capacity, it has been the primary focus of prior research. Crucially, we argue that the final term, the predictive variance, is largely overlooked in the design of deterministic neural forecasters. By default, these models yield a single output ($c=1$), leaving this error component from the inherent uncertainty of a neural network unresolved. Unlike classical methods often constrained by closed-form derivations, NFs possess the intrinsic flexibility to generate multiple outputs, thereby actively reducing this variance.
@@
Post combination of multiple forecasts. Given a set of candidate outputs from an NFM, there exist aggregation strategies to synthesize a single, integrated forecast that is more accurate than any individual candidate (combiningreview). We represent this aggregation as a functional $g_{c}$ that produces the post-combined forecast:
@@
In Sec. 3.3, we empirically demonstrate that the performance of an ’oracle’ combination on real-world datasets is consistent with the hypothesis. However, constructing an optimal combination in practice is non-trivial (combiningreview) since intricate combination methods practically underperform the simple average (combinationpuzzel). Consequently, a primary objective of our proposed framework is to bypass explicit combination design and instead enable the forecasting paradigm to implicitly learn itself how to leverage the information contained within these multiple forecasts.
@@ 2.2 Novel Forecasting Paradigm
The current frameworks in LTSF involve two main strategies: 1-step Auto Regressive (AR(1)) and Direct Output (DO). While AR models perform consistently with recurrent state space models like RNNs (lstm4ltsf), they are known to suffer from significant error accumulation in long-horizon forecasting and are empirically evidenced to largely fall behind DO in LTSF (zhou2021informer), making a strong reason for the dominant position of DO in LTSF. However, the shortcomings of the DO approach itself have been largely overlooked. We argue that despite its simplicity, the most significant weakness of DO, particularly when compared to AR, is its lack of inherent causality in the structure of the predictive representations.
To illustrate this, we conduct a simple experiment in Fig.3: we consecutively split a forecast of length $L$ into $m$ non-overlapped chunks with length $L/m$, and each is predicted by an independent linear head from a shared representation. The final forecast is a concatenation of these segments. Empirically, it turns out that the performance of this multi-headed model is nearly identical to a standard DO model with a single head predicting the full horizon. This result suggests that the NF is almost unaware of the temporal relationships within the future sequence and makes the prediction without considering the interval between the input history and the output future. As a result, it learns to forecast each segment independently, which seems anti-intuitive.
Figure 3: We make the forecast by applying independent heads on several non-overlapped chunks. The right figure shows the training loss curve in different settings.
According to this observation, and combining our heuristic idea from VRH, we propose a new forecasting paradigm as follows,
Definition 1 (Boosted Direct Output (BDO)).
@@
Remark. BDO generalizes DO but is distinct from AR. Unlike existing AR paradigms (deepar) which feed predictions back to yield the next step consecutively, BDO regenerates the horizon from the start point at each iteration. Thus, every intermediate output is a complete, independent sub-forecast rather than a dependent fragment of a single trajectory.
Intuitively, BDO recursively generates forecasts for progressively longer horizons, reusing predictions from previous stages. This incorporates an AR-like causal structure into the forecasting process, while retaining the patch-wise output characteristic of DO that mitigates significant error accumulation. A balance between these two properties can be achieved by properly setting the number of recursive stages, $N$ (see Appendix B.1 for theoretical analysis). Furthermore, since short-term forecasting is generally an easier task than long-term forecasting, BDO effectively creates a learning curriculum and tends to build hierarchical representations as illustrated in Fig.2. We can reasonably expect this paradigm to outperform both pure AR and pure DO strategies, especially in the challenging LTSF setting.
@@ 2.3 Model Architecture
Figure 4: Model Structure of ReNF. PE denotes optional embeddings, and the $\beta$ version adds additional skip-connections for deeper representations.
We construct our NFs using only MLP and linear layers for: 1) As foundational deep learning modules, improvements demonstrated on them are broadly applicable and convincing. 2) Recent work has shown that MLP-based architectures perform well for a wide range of forecasting tasks (ekambaram2023tsmixer).
Fundamentally, we build two NF variants, ReNF-$\alpha$ and ReNF-$\beta$, which differ in their degree of non-linearity for handling datasets of varying complexity. The overall architecture is shown in Fig.4, which is characterized by two main principles.
Block-wise Supervision. To adopt the BDO strategy, we stack multiple blocks, and each consists of a linear layer for representation projection, followed by an MLP for nonlinear transformation. Particularly, the end of each block is equipped with a dedicated linear head that maps the internal representation back to the data space, steering it to function as a sub-forecaster for horizon $h_{j}$.
Generalized Forecast Combination. In the recursive BDO process, each sub-forecast is concatenated with the original input to form the new input for the next block. This operation is trivial yet crucial in our design as it allows an implicit generalized forecast combination as follows,
[…]
where we characterize the $n$-th MLP block as an operator composed by $f:\mathbb{R}^{h}\rightarrow\mathbb{R}^{h_{n}}$ and $\varphi:\mathbb{R}^{h_{n-1}}\rightarrow\mathbb{R}^{h_{n}}$. Thus, if we densely connect all outputs, every intermediate stage of the NF will recursively learn the implicit combination and finally yield a complex composed forecast.
We employ RevIN (kim2021reversible) as the pre-normalization for the initial input data to reduce the distribution discrepancy between the training and evaluation phases. For consistency, we also apply pre-normalizations to the input of each sub-forecaster. Additionally, we apply dropout before the initial linear projection to prevent the model from being overly dependent on the history observation and the concatenated information in the data space. To allow for representation flow with diverse depths, the ReNF-$\beta$ variant also incorporates skip-connections between the representation spaces.